OpenStudy (anonymous):

Solve 2(x + 1) = 2x + 2. All real numbers No solution 0 1

OpenStudy (solomonzelman):

\(\normalsize\color{black}{ 2(x + 1) = 2x + 2 }\) factor out of 2 on the right. OR INSTEAD, expand the left side, \(\normalsize\color{black}{ 2x +2= 2x + 2 }\) Either way:)

OpenStudy (solomonzelman):

This is what you get after factoring the right side out of 2. \(\normalsize\color{black}{ 2(x +1)= 2(x + 1) }\)

OpenStudy (anonymous):

A) all real numbers because 2(x+1)=2x+2= 2x+2=2x+2 since its 2 things equal to each other that are the same thing its all real number. any solution that's a real number works

OpenStudy (anonymous):

ty lacross

OpenStudy (anonymous):

np :)